The cost to produce Part A was $20 per unit in 2019. During 2020, it has increased to $22 per unit. In 2020, Pharoah Company has offered to supply Part A for $15 per unit. For the make-or-buy decision,

incremental costs are $2 per unit.
differential costs are $7 per unit.
net relevant costs are $2 per unit.
incremental revenues are $7 per unit.

Oriole Company is considering the following alternatives:

Alternative A Alternative B
Revenues $38000 $54000
Variable costs 22800 22800
Fixed costs 10000 16000


What is the incremental profit?

$6000
$10000
$5200
$0
Answer #1

1. Differential costs = $22-$15 = $7 per unit

differential costs are $7 per unit is correct answer.

2. Profit on A = $38000-22800-10000 = $5,200

Profit on B = $54000-22800-16000 = $15,200

Incremental profit = $15200-5200 = $10,000